OpenStudy (anonymous):

. The new releases make up 1/5of the available DVDs at a DVD store. 1/5d=64 The equation below is used to determine the total number of DVDs (d) in the store. What is the total number of DVDs in the store? 69 320 325 13

OpenStudy (the_fizicx99):

What equation? lol

OpenStudy (anonymous):

sorry typed it wrong

OpenStudy (anonymous):

The new releases make up 1/5of the available DVDs at a DVD storeThe equation below is used to determine the total number of DVDs (d) in the store. 1/5d=64 What is the total number of DVDs in the store? 69 320 325 13

OpenStudy (the_fizicx99):

I'd say D, 1/5d * 1/5 = 64*1.5 = 12.8 -> rounded would be 13 :3 sweg sweg

OpenStudy (anonymous):

its b but thanks anyway :)

OpenStudy (the_fizicx99):

Oh, 64*5 / 1 = 320.. sowwie.

OpenStudy (anonymous):

no your fine thanks for the help :)